A new virus is released on the internet; the administrator of a department's Local Area Network(LAN)is given 5 minutes by a manager to estimate the impact. The administrator samples 12 of the PC's connected to the LAN, and finds that 7 are infected. Use proportion to estimate the number of infected PC's if there are a total of 117 PC's connected to the LAN.
Answer by stanbon(75887)

7/12 = x/117
x = (7/12)*117 = 68.25
Round down to x = 68 of the 117 are infected.
=================
Cheers,
Stan H.